An undergraduate student weighed 20g of sodium carbonate salt. If Na=23, O=16, carbon = 12and H=1, what is the mole of the hydroxide?
1
Expert's answer
2015-10-21T06:02:16-0400
Sodium carbonate has a formula Na2CO3.
Number of moles of sodium carbonate is:
n = m/M = 20/(23*2+12+16*3) = 0.19mol

Answer: 0.19mol